# San Jose/Seattle, US https://bytedance.github.io/vidi-website/

### v1

@@ 4.2 Temporal Retrieval
Table 8: Video QA performance of Vidi2 on popular benchmarks.
We also present the ROC curve for different metrics in Figure 5, and we can observe that Vidi2 consistently outperforms all competitors in IOU, precision, and recall across the entire range of thresholds. It shows Vidi2’s strength in fine-grained temporal understanding, a critical capability for real-world video editing tasks.
4.3 Generic Video QA
As a video understanding foundation model, Vidi2 also supports video question answering, which is a major improvement over the previous Vidi model. We evaluate Vidi2 and several popular multimodal models on three representative video QA benchmarks: LVBench [wang2025lvbench], LongVideoBench [wu2024longvideobench], and VideoMME [fu2025video]. While Vidi2 still lags behind the leading proprietary model, Gemini-2.5-Pro [comanici2025gemini], in general video QA performance, it achieves competitive results compared to the most popular open-source multimodal model of similar scale, namely Qwen2.5-VL-7B [qwen2.5-VL]. These results demonstrate that Vidi2 possesses solid multimodal reasoning ability despite being primarily optimized for retrieval and grounding tasks.
@@ 4.3 Generic Video QA
Figure 7: Example of plot understanding application.
5 Appplications
In view of Vidi2’s strong capabilities in temporal retrieval, spatio-temporal grounding, and generic video question answering, it is well suited to support a wide range of video editing tasks, including highlight extraction, plot understanding, and storyline-based video creation.
5.1 Highlight with Title
@@
The strong temporal modeling capability of Vidi2 enables straightforward adaptation to downstream tasks such as highlight prediction. Given a long raw video, the model can automatically generate multiple highlight segments along with concise titles, allowing users to directly publish the resulting short clips. As shown in Figure 6, the generated titles align closely with the corresponding video content, demonstrating both temporal precision and semantic understanding.
5.2 Plot Understanding
Popular movies and TV series often feature complex relationships among multiple characters. Performing character- or plot-based video cutting is therefore time-consuming, even for professional editors. To enable automatic editing in such contexts, a model must comprehend character identity, interactions, and event causality—for example, understanding that “one character raises an arm to hit another but misses and strikes a friend instead.” Such scenarios require robust spatio-temporal perception, character recognition, commonsense reasoning, and background knowledge. As shown in Figure 7, Vidi2 supports plot-level understanding and reasoning through a dedicated pipeline, demonstrating its potential in real-world storytelling applications.
5.3 Storyline-based Video Creation
Creating an engaging short video is akin to producing a miniature film with a coherent narrative. A high-quality video depends not only on appealing content but also on an effective storyline and professional-level editing composition. Next-generation video creation systems should act like professional editors, i.e., cutting clips according to a narrative structure while considering package elements such as effects, transitions, music, and stickers. Vidi2 demonstrates the ability to generate end-to-end video creation instructions conditioned on multiple input videos. As shown in Figure 8, the model takes six videos as input and outputs a complete instruction set for producing a publishable, storyline-driven video. The final rendered output includes narration, music, animations, and transitions, showcasing the model’s potential to automate the entire creative editing workflow.
Figure 8: Example of storyline-based video creation application.
@@ 6 Conclusion
In this report, we present Vidi2, which achieves unprecedented end-to-end spatio-temporal grounding performance while maintaining strong video question answering capabilities. It serves as a foundation model for next-generation video creation and understanding scenarios. To evaluate models in realistic video editing contexts, we introduce VUE-STG, a comprehensive benchmark that advances prior spatio-temporal grounding datasets in four aspects: video duration, query format, annotation quality, and evaluation metrics. We also upgrade our previous temporal retrieval benchmark to VUE-TR-V2, featuring a more balanced video-length distribution (including more long and ultra-long videos) and more realistic, user-style queries. In addition, we evaluate Vidi2 on public video QA datasets to assess its general multimodal reasoning ability. Notably, Vidi2 significantly outperforms proprietary systems such as Gemini 3 Pro (Preview) and GPT-5 on both VUE-STG and VUE-TR-V2, demonstrating state-of-the-art spatio-temporal understanding. It also performs on par with the leading open-source video LLM of similar scale, i.e., Qwen2.5-VL-7B, on video QA. Finally, we outline three promising real-world applications that Vidi2 could enable in future demonstrations. Together, these results highlight the model’s strong spatio-temporal understanding and multimodal reasoning capabilities, forming a robust foundation for intelligent, human-like video editing and creation systems.
@@ 7 Contributors
Core Contributors - Research (alphabetical order)
Chia-Wen Kuo, Dawei Du, Fan Chen, Guang Chen, Sijie Zhu, Wen Zhong, Xin Gu, Zhenfang Chen.
@@
Celong Liu, Haojun Zhao, Tong Jin.
@@
Longyin Wen, Xiaohui Shen.
@@
Chuang Huang, Haoji Zhang, Lingxi Zhang, Lu Guo, Lusha Li, Qihang Fan, Qingyu Chen, Rachel Deng, Stuart Siew, Weiyan Tao, Zuhua Lin.

### v2

@@ 4.3 Generic Video QA
Vidi2.5 45.2 58.9 63.6
Table 9: Video QA performance on popular benchmarks.
@@
Figure 9: Example of chapter summary application.
@@
The strong temporal modeling capability of Vidi2 enables straightforward adaptation to downstream tasks such as highlight prediction. Given a long raw video, the model can automatically generate multiple highlight segments along with concise titles, allowing users to directly publish the resulting short clips. As shown in Figure 8, the generated titles align closely with the corresponding video content, demonstrating both temporal precision and semantic understanding. Similarly, as shown in Figure 9, the model can generate chapters with summary and titles for a long raw video.
4.4 Plot Understanding and Reasoning
Inference Setup for LMMs. For the Character Track of VUE-PLOT, we design task-aligned prompts and structured output formats to encourage the evaluated MLLMs to generate stable and parsable outputs when performing active speaker transcription and face tracking. This design facilitates a fair comparison with Vidi2.5. In the following, we detail the input and output configurations for the compared models.
Gemini 3 and Qwen3-Omni Input/Output Format: We evaluate Gemini 3 Pro Preview [2] and Qwen3-Omni-Thinking [22], leveraging their multimodal capabilities to process both video and audio inputs. For all experiments, the video inputs are sampled at $1$ FPS. To minimize the influence of prompt variations on the evaluation results, we apply the same structured prompt to both models and enforce a unified output format via a JSON schema at the API level. Notably, the structured output functionality of Qwen3-Omni [22] is supported in the latest version of vLLM******https://github.com/vllm-project/vllm. Under these constraints, both models are required to produce a JSON object containing precise timestamps and normalized coordinates, i.e.,
Transcribe the active speaker’s speech and track their face
using bounding boxes. Output only a single JSON object and
do not output explanation.
Timestamp format: Total seconds (float).
Bounding box format: [x0, y0, x1, y1] with values in [0, 1] normalized
to the video frame, tightly enclosing the speaker’s face.
Sampling strategy: Sample at the first visible frame and every subsequent
integer second.
Example:
  "text": "Hello everyone.",
  "start": 62.4,
  "end": 65.0,
  "boxes": [
    {"timestamp": 62.4, "box_2d": [0.400, 0.150, 0.600, 0.350]},
    {"timestamp": 63.0, "box_2d": [0.405, 0.155, 0.605, 0.355]}
Track Category Metric (%) Models
Vidi 2.5 Gemini 3 Pro GPT 5 Qwen3 VL Qwen3 Omni
Char.                - tIoU 71.63 66.04 - - 50.68
sIoU 55.89 13.24 - - 6.80
WER $\downarrow$ 23.20 29.00 - - 58.12
Rea. Perception and Understanding Acc 66.25 80.42 55.83 35.00 40.83
Speech, Audio, and Sound Effect Reasoning Acc 74.43 66.03 55.34 45.42 27.10
Social Cognition and Knowledge Integration Acc 62.78 62.15 53.94 31.23 23.03
Narrative and Structural Understanding Acc 55.17 59.48 51.72 25.86 25.00
Professional Filming and Editing Techniques Acc 61.35 50.92 55.21 30.67 24.54
Overall Acc 64.33 64.58 54.37 33.94 28.01
Table 10: Evaluation results on the VUE-CHAR benchmark, reporting temporal intersection-over-union (tIoU), spatial intersection-over-union (sIoU) and ASR errors in the Character (Char.) and question accuracy in the Reasoning (Rea.) track.
Performance Evaluation.
We present the detailed results in terms of grounding precision and reasoning accuracy in Table 10. Notably, Vidi2.5-Think significantly outperforms Gemini 3 Pro Preview [2] for spatial IoU on VUE-CHAR ($42\%$) and achieves the lowest word error rate. For detailed analysis on the character track, although Gemini 3 Pro Preview maintains a slight edge in overall accuracy due to strong base perception, Vidi2.5-Think excels on Speech, Audio, and Sound Effect Reasoning ($74.43\%$ vs $66.03\%$) and Professional Filming and Editing Techniques ($61.35\%$ vs $50.92\%$). We can conclude that the superior audio-visual alignment of Vidi2.5-Think ensures a better experience for understanding complex cinematic narratives and subtle character interactions. We show Exemplar cases of Vidi2.5-Think in Figure 10 and Figure 11. Vidi2.5-Think is able to densely localize the active speakers and recognize what they are talking about (see Figure 10). Vidi2.5-Think also provides step-by-step reasoning to arrive at the correct answer, explicitly decomposing the complex plot understanding process into a sequence of interpretable intermediate inferences that mirror human-like temporal and causal reasoning over video content (see Figure 11).
Figure 10: Exemplar cases of Vidi2.5-Think for dense speaker localization and speech recognition in videos. The ground-truth annotation and the model prediction are are marked with green and red colors in the input video for better visualization.
Figure 11: Exemplar cases of Vidi2.5-Think for plot reasoning in videos. Vidi2.5-Think provides step-by-step reasoning to arrive at the correct answer. The correct option is marked with green color for better visualization.
5 Applications
Beyond standard multimodal understanding benchmarks, Vidi2.5 can be adapted through post-training to support a broad range of downstream applications that require structured reasoning, long-horizon planning, and coordinated multimodal decision-making. In this work, we present video editing planning as a representative case study to demonstrate these capabilities in a concrete and practical scenario.
5.1 Video Editing Planning
Video editing planning aims to determine a set of interdependent high-level editing decisions that collectively define how a final video is created from raw assets. Given a collection of raw assets (e.g., images and videos) and optional user prompts (i.e., user-specified editing requirements), the objective is to generate an editing plan that specifies the narrative structure, narration content, audio attributes, and visual editing intent.
In our setting, the editing plan is represented as a high-level textual or structured representation that unifies narrative structure, audio design, and visual intent. This representation enables downstream systems to perform audio synthesis, visual effect parameterization, and video rendering in a coordinated manner.
5.2 Editing Plan Specification
The editing plan serves as a semantic blueprint, defining a set of high-level decisions that together specify the structure, presentation, and intent of the final video. As shown in Figure 12, these decisions act as a unified directive for downstream execution systems.
Narrative structure
defines the hierarchical and temporal orchestration of the video, encompassing clip selection, segment extraction, and segment ordering. It establishes the temporal and semantic backbone of the edit and governs how visual content is arranged to convey the intended narrative.
Voiceover content
specifies the narrated material aligned with the narrative structure, including semantic content, delivery style, and temporal alignment. It determines how narration supports the storyline and provides guidance for downstream speech synthesis.
Audio attributes
characterize auditory elements such as music style, mood, rhythmic profile, and speaker attributes, which contextualize and complement the visual narrative. This attribute set establishes the auditory framing of the video and promotes coherence between sound design and visual progression.
Visual editing intent
articulates stylistic and rhythmic presentation directives in semantic terms, including transition design, emphasis cues, and stylistic directives. This specification communicates how visual effects should be applied to support narrative structure and pacing.
Together, these decision categories define the specification space of the editing plan, capturing the intended structure and presentation of the video while remaining agnostic to execution-specific details.
Figure 12: Overview of the Vidi-Edit execution pipeline. The generated editing plan specifies high-level semantic intent, including narrative structure, voiceover content, audio attributes, and visual editing directives, which is subsequently translated into concrete operations, such as music recommendation, speech synthesis, visual effect retrieval, and video rendering. This design explicitly separates high-level planning from execution, allowing the model to focus on semantic reasoning while downstream systems handle implementation.
5.3 Post-Training for Editing Plan Generation
To support video editing planning, we post-train Vidi2.5 to align its multimodal reasoning capabilities with the structured requirements of editing plan generation. We refer to this post-trained configuration as Vidi-Edit. This post-training stage does not modify the underlying model architecture; instead, it adapts the pre-trained representations to produce editing plans that conform to the specification described above.
Post-training leverages additional task-specific supervision in the form of multimodal inputs paired with textual or structured editing plans. The supervision signal emphasizes adherence to the planning format and semantic consistency across narrative structure, voiceover content, audio attributes, and visual editing intent, enabling the model to express its reasoning outcomes in a coherent planning-oriented representation.
5.4 Editing Plan Implementation
The editing plans generated by Vidi-Edit are realized through a structured execution pipeline that translates high-level planning intent into concrete editing actions. As shown in Figure 12, the generated plans, expressed in natural language or structured text, are first processed by a translation stage that maps semantic specifications to executable components.
During this interpretation stage, different aspects of the editing plan are grounded through specialized modules. Music attributes specified in the plan are retrieved from a curated music database, while the narration content is synthesized by a text-to-speech (TTS) generator according to the speaker attributes. The visual editing intent is mapped to the corresponding visual effect primitives through a dedicated effect retrieval and parameterization module.
The resulting audio assets and visual effect parameters are then passed to an in-house rendering system, which integrates video clips, synthesized audio, textual overlays, and visual effects to produce the final high-quality videos. This modular execution pipeline cleanly separates high-level editing decisions from their realization, allowing the planning model to focus on semantic reasoning while execution components handle asset selection, synthesis, and rendering.
Under this design, the planning-and-execution pipeline can be further extended into an editing agent architecture, where the Vidi-Edit serves as a high-level decision-making module and the execution pipeline grounds these decisions into concrete editing actions.
5.5 Qualitative Examples
We present qualitative examples to examine the editing plans produced by Vidi-Edit and to illustrate how high-level editing intent is reflected in the resulting outputs. Figures 13 14 15 16 show representative rendered video frames from which narrative structure, clip arrangement, and voiceover progression can be inferred.
In these examples, Vidi-Edit generates coherent editing plans that organize raw video clips into an explicit temporal structure. The clips are arranged according to the planned storyline, including context setup (e.g., an opening hook), content development, and a concluding wrap-up. Voiceover content is aligned with this temporal structure, providing a consistent narrative thread that follows the planned progression of the video.
While the figures visualize clip ordering and voiceover progression, other planning attributes, such as audio design and visual effect specifications, are not directly observable in static frames. To complement these visualizations, additional qualitative results and example videos are provided on the project website, along with an interactive interface for exploring the model’s capabilities.
Figure 13: Qualitative example of an editing plan generated by Vidi-Edit for a single-topic scenario (zoo animals). All input assets are captured within the same zoo environment. The model identifies salient moments across multiple animal clips and arranges them into a coherent narrative with implicit temporal structure, rather than merely concatenating visually similar footage.
Figure 14: Qualitative example of an editing plan generated by Vidi-Edit for a single-topic urban travel scenario. Given heterogeneous user assets from a city visit (e.g., transportation, streets, nightlife, and social moments), the model constructs a unified storyline that captures temporal progression and experiential flow across diverse scenes.
Figure 15: Qualitative example of an editing plan generated by Vidi-Edit for a single-topic urban scenario set in Seattle. Each input consists of only a few seconds, providing limited temporal context in isolation. Despite this constraint, the model infers high-level semantic relationships across clips and organizes them into a coherent narrative, demonstrating robustness to sparse and fragmented inputs.
Figure 16: Qualitative example of an editing plan generated by Vidi-Edit for a mixed-content scenario. The input assets are semantically diverse and lack an obvious temporal or thematic order. Vidi-Edit selectively identifies and arranges appropriate clips to construct a coherent storyline, demonstrating its ability to perform high-level narrative planning beyond low-level visual similarity or naive clip concatenation.
@@ 6 Conclusion
In this report, we first present Vidi2, which achieves unprecedented end-to-end spatio-temporal grounding performance while maintaining strong video question answering capabilities. It serves as a foundation model for next-generation video creation and understanding scenarios. To evaluate models in realistic video editing contexts, we introduce VUE-STG, a comprehensive benchmark that advances prior spatio-temporal grounding datasets in four aspects: video duration, query format, annotation quality, and evaluation metrics. We also upgrade our previous temporal retrieval benchmark to VUE-TR-V2, featuring a more balanced video-length distribution (including more long and ultra-long videos) and more realistic, user-style queries. In addition, we evaluate Vidi2 on public video QA datasets to assess its general multimodal reasoning ability. Notably, Vidi2 significantly outperforms proprietary systems such as Gemini 3 Pro Preview and GPT-5 on both VUE-STG and VUE-TR-V2, demonstrating state-of-the-art spatio-temporal understanding. It also performs on par with the leading open-source video LLM of similar scale, i.e., Qwen2.5-VL-7B, on video QA. Finally, we outline three promising real-world applications that Vidi2 could enable in future demonstrations. Together, these results highlight the model’s strong spatio-temporal understanding and multimodal reasoning capabilities, forming a robust foundation for intelligent, human-like video editing and creation systems.
In addition, we introduce the latest Vidi2.5 that adopts RL training to consistently improve the performance of TR, STG and video QA over the Vidi2 base model. We also introduce a thinking model, Vidi2.5-Think, with inference-time scaling to handle complex reasoning scenarios in video creation. Specifically, we focus on plot understanding in this report and propose the VUE-PLOT benchmark to comprehensively evaluate the fine-grained perception and complex reasoning capability in this task. Vidi2.5-Think outperforms Gemini 3 Pro Preview on fine-grained character understanding by a large margin, and achieves comparable performance on complex plot reasoning tasks.
Beyond benchmark-driven evaluation, we further demonstrate that Vidi2.5 can be adapted through post-training to support practical video creation workflows. Using video editing planning as a representative application, the post-trained variant Vidi-Edit generates structured editing plans that specify narrative structure, audio attributes, and visual editing intent from raw video assets. This application illustrates how the spatio-temporal understanding and multimodal reasoning capabilities of Vidi2.5 can be translated into executable planning representations for real-world video editing systems.
@@ 7 Contributors
Core Contributors - Research - Foundation (alphabetical order)
Chia-Wen Kuo, Dawei Du, Fan Chen, Guang Chen, Haoji Zhang, Sijie Zhu, Xin Gu, Zhenfang Chen.
Core Contributors - Research - Application (alphabetical order)
Feng Gao, Lu Xu, Wen Zhong, Ye Yuan, Yiming Cui.
@@
Fanding Lei, Haojun Zhao, Tong Jin.
@@
Longyin Wen [^1] Corresponding author., Xiaohui Shen.
@@
Chuang Huang, Jin Liu, Jingjing Zhuge, Lili Fang, Lingxi Zhang, Lu Guo, Lusha Li, Qihang Fan, Rachel Deng, Shaobo Fang, Shu Zhang, Stuart Siew, Weiyan Tao, Yicheng He, Zhihua Wu, Zuhua Lin.